NA4:

ACTIVITY: SOLUTION CONCENTRATION VS. CONDUCTIVITY A chemistry student carried out an experiment with a conducting apparatus (ammeter) similar to the one below. This ammeter measured in milliamperes (mA). The following data was taken. Solution Reading 0.1 M H2SO4 150 mA 0.1 M Ba(OH)2 150 mA To 30 mL of the Ba(OH)2 solution, 10 mL portions of H2SO4 were added until a total of 50 mL of H2SO4 were used. The following results were recorded. DATA TABLE Total H2SO4 Added Meter Reading Observations 0 mL 150 mA Ba(OH)2 and H2SO4 clear, colorless 10 mL 65 mA milky white precipitate forms 20 mL 31 mA more precipitate forms 30 mL 0 mA precipitate heavy and settles 40 mL 29 mA no added precipitate seen to form 50 mL 62 mA no change seen How does this evidence indicate that the reaction has occurred between ions? Please Help with questions 3 and 4.

NA4:

Does the conductivity increase or decrease? Does the number of ions in solution increase, decrease, or remain constant? What is the indicator of the number of ions in solution?

NA4:

Why does it conduct more before and after this minimum point? Please Help

justjm:

Q1. Form your double replacement equation \(Ba\left(OH\right)_{2}\ +\ H_{2}SO_{4}\left(aq\right)\ →\ Ba_{2}SO_{4}\left(s\right)+H_{2}O(l)\) Note: this is unbalanced Ba2SO4 is insoluble because sulfate is only soluble with NH4+ and IA metals. Thus a precipitate of Ba2SO4 forms. This is a precipitation rxn, and yes a reaction has occurred. [H+] and [OH-] would remain the same to form H2O, but [Ba+] and [SO4 2-] would decrease. Q2. Conductivity will decrease because as more precipitate forms, the Ba+ and SO4 2- ions don't dissociate, and in order to be an electrolyte, the ions must dissociate. Q3. sORRY not sure, I think it's just the mA Q4. Sorry I'm not sure

NA4:

Thank you justjm so much :)

justjm:

np, I think @Vocaloid might assist you better

NA4:

Ok, thank you

justus:

Actually justjm is better

justus:

assistance

NA4:

Ok, I honestly just need help with the last two questions which are Does the number of ions in solution increase, decrease, or remain constant? What is the indicator of the number of ions in solution? Why does it conduct more before and after this minimum point?

justus:

I can't answer

NA4:

It's ok :)

justus:

Good luck.

NA4:

Thank you :)

justjm:

didn't see the table actually, I can answer 2 Does the number of ions in solution increase, decrease, or remain constant? It increases after 30 mL of H2SO4 is added bc no precipitate is showing up and mA increases based on the table. What is the indicator of the number of ions in solution? It's mA. Why does it conduct more before and after this minimum point? --> this I do not know, sorry.

NA4:

It's ok, I do appreciate the help though.